MICQ(1) MICQ(1) mICQ−ICQmicq[-h][-c][-b<>][-i<>][-v[]][-l<>][-u<UIN>][-p<>][-s<>][-C<>] mICQICQ−,ICQ,Mirabilis.mICQ,,,,SMS,,,,−,UTF−8,,. −h|−−help . −c|−−nocolor . −b|−−basedir<> ,mICQ,micqrc(5). −i|−−i18n<> . −v|−−verbose[] . −l|−−logplace<>|<> . −u|−−uin<uin> uin −p|−−passwd<> -u. −s|−−status<> ,inv,dnd,occ,na,away,ffc,off[line]. −C|−−cmd<> mICQeg. ()mICQUIN.. msg(<tab>msg,),addupdate,other,andabout−ICQ,,help.micq(7). mICQLinux,DOS,OS/2,Win32,BeOS,AmigaOS,GNUHURDunix−.mICQ0.5 Linux,AIX,HPUX,Tru64,AmigaOS,BeOSunix−.,,. 161998 .micq.12251998. 052001 ..,mICQBSDmICQ. 102002 0.4.9GPL,−ICQ8,Ru¨ digerKuhlmann. mICQmicq-list@micq.org.http://www.micq.org/pipermail/micq-list/. HTTPhttp://www.micq.org/download. CVS: cvs−d:pserver:anonymous@micq.org:/var/lib/cvscomicq mICQ 1 MICQ(1) MICQ(1) . micqrc(5),micq(7) ˜ /.micqrc mICQ−. ˜ /.micq/ ,mICQ(BASE). BASE/micqrc mICQ.. BASE/status mICQ.. BASE/i18n/*.i18n mICQ. BASE/contacts/ ,-. BASE/history/ . BASE/files/ ,. BASE/remote-control FIFO,.,. BASE/scripting FIFO,. PREFIX/share/micq/*.i18n ,mICQ. Ru¨ digerKuhlmann<micq@ruediger-kuhlmann.de>(). <ja2morrison@student.math.uwaterloo.ca>,<jordi@sindominio.net>DebianGNU/Linux <micq@ruediger-kuhlmann.de>. mICQ 2